[Brahms-l] Shift report 20050527 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri May 27 2005 - 07:52:18 EDT
Shifter leader was rk
Shifters were 


1. Problem with RF that caused debunching of blue beam turned up to be more serious. Had beam for less than half hour.
2. One physics run - 14981. Pedestals.
3. BB3 Mainframe has problems again. Since moving CC counters to other mainframes it affects mainly BB large tubes - probably some more effort should be put to make the mainframe stable and working.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri May 27 07:52:32 2005

This archive was generated by hypermail 2.1.8 : Fri May 27 2005 - 07:52:53 EDT